zarasht:

A sample of argon has a volume of 1.2 L at STP. If the temperature is increased to 22°C and the pressure is lowered to 0.80 atm, what will the new volume be, in L?

justjm:

Use combined gas law \( \frac{P_1V_1}{T_1} = \frac{P_2V_T}{T_2}\) \(V_1=1.2L\) \(T_1=273.15K\) \(P_1=1atm\) \(T_2=(22°C+273.15)K\) \(P_2=0.80atm\) Solve for \(V_2\) by either rearranging the formula or solving for it using algebra.